OpenStudy (princesssleelee):

Choose one of the factors of x15 − 64. x5 − 4 x10 − 4x5 + 16 x10 − 4 x15 + 8x + 16

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle x^{15}-64 }\) ?

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle (x^{5})^3-4^3 }\)

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle b^3-a^3=(b-a)(b^2+ba+a^2) }\) (this is the rule)

OpenStudy (princesssleelee):

I am thinking A from the info you have given so far.

OpenStudy (solomonzelman):

Yes, that was quick:)

OpenStudy (solomonzelman):

is this question "for all that applies" ?

OpenStudy (solomonzelman):

I just want to say, by asking this question, that there is another correct answer.

OpenStudy (princesssleelee):

I got a 100%, thank you again!

OpenStudy (mathstudent55):

There is no other correct choice. A was the only one.